OpenStudy (anonymous):

4(x + 1) = 4x + 10 0 6 Any Real Number No Solution

OpenStudy (anonymous):

4(x+1)=4x+10 4x+4 = 4x+10 4=10? Not possible, no solution

OpenStudy (goformit100):

No Solution

OpenStudy (goformit100):

as 4x gets cancels out